Heads of lettuce that are normally priced at $0.78 each are now being offered for a special price of 3 for $2. If Mr. Graboski wants to buy 6 heads of lettuce, how much will he save if he buys them at the sale price instead of the regular price?
1
Expert's answer
2012-09-04T10:42:09-0400
If Mr. Graboski buys 6 heads of lettuce at the regular price,
he pays 6 *
0.78 = 4.68$.
But if a special price is proposed, he pays $2 for each 3
heads,
so he spends 2 * 2 = 4$ as he needs 6 heads.
So, Mr. Graboski saves
4.68 - 4 = 0.68$.